OpenStudy (anonymous):

find the sum of 35 term of an AP whose "P"th term is (p/7) + 12

OpenStudy (anonymous):

is the answer 425?? if it is correct i will elaborate

OpenStudy (anonymous):

no

OpenStudy (mathmagician):

S=35*12+1/7(1+2+3+...+35)=420+1/7*(1+35+2+34+...+(17+19)+18)=320+1/7*(36*17+18)=320+90=410

OpenStudy (anonymous):

no

OpenStudy (anonymous):

answer is 510

OpenStudy (anonymous):

(p+84)/7 now apply the summation rule to the above with upper limit 35 and lower limit 1 \[((\sum_{1}^{35}p + \sum_{1}^{35}84))/7\] =(35*18 + 84*35)/7 =510

OpenStudy (mathmagician):

oh, really- 12*35=420, not 320, mind calculation error :D so the answer is 510

OpenStudy (anonymous):

did you get my method???@srinidhija

OpenStudy (anonymous):

yeah , corrct sinc pth term is (p/7) +12 => 1st term = 1/7 + 12= 85/7 2nd trm= 86/7 ==> common difference= 1/7 => sum of 35 term = 35/2 (2x85/7 + (35-1)1/7) =5 (85 +17) =510

OpenStudy (anonymous):

if you use the principle of summation you sum becomes shorter and you can solve them faster!!!! nevermind absorb the method which suits you the most